### v1

Abstract Sequence alignment is a cornerstone technique in computational biology for assessing similarities and differences among biological sequences. A key variant, sequence-to-graph alignment, plays a crucial role in effectively capturing genetic variations. In this work, we introduce two novel formulations within this framework: the Gap-sensitive Co-Linear Chaining (Gap-CLC) problem and the Co-Linear Chaining with Errors based on Edit Distance (Edit-CLC) problem, and we investigate their computational complexity. We show that solving the Gap-CLC problem in sub-quadratic time is highly unlikely unless the Strong Exponential Time Hypothesis (SETH) fails—even when restricted to binary alphabets. Furthermore, we establish that the Edit-CLC problem is NP-hard in the presence of errors within the graph. These findings emphasize that incorporating ”co-linear” structures into sequence-to-graph alignment models fails to reduce computational complexity, highlighting that these models remain at least as computationally challenging to solve as those lacking such prior information.
Keywords Co-linear chaining; Sequence alignment; Algorithm; Complexity
MR(2020) Subject Classification 05C85; 68Q17; 68R10; 92D20; 92-08
1. Introduction
Genomic mutations shape the genomes of populations, resulting in genetic divergence among individuals. Consequently, the comparative analysis of genetic sequences enables the identification of genetic similarities and differences [22], offers insights into the origin, evolution, and adaptability of species [27], and elucidates the evolutionary relationships and functional dynamics of biological sequences [8]. Sequence alignment is widely employed in the analysis of DNA, RNA, and protein sequences, serving as a crucial method for exploring gene families, protein structures, and functional characteristics [22].
Sequence alignment aims to identify the optimal correspondence among two or more sequences, revealing their similarity levels and specific variations. By incorporating matches, substitutions, and gaps or insertions, sequences are adjusted to maximize alignment quality according to predefined criteria [5].
Biologists and computer scientists have devised a variety of algorithms to tackle sequence alignment, a fundamental task in bioinformatics. Among the most classical approaches are global and local alignment methods [4]. Global alignment seeks to maximize similarity across the entire length of the sequences, providing a comprehensive comparison, whereas local alignment identifies regions of high similarity within subsequences, even if the rest of the sequences differ significantly.
Given the vast amount of sequence data and the need to evaluate numerous possible alignments, sequence alignment is inherently computationally demanding. To address this challenge, researchers have introduced efficient optimization techniques and heuristic strategies. These include dynamic programming [29], which guarantees optimal solutions by breaking the problem into smaller subproblems, the widely used BLAST (Basic Local Alignment Search Tool) [16] that accelerates searches through heuristic filtering, and the sensitive Smith-Waterman algorithm [28] for precise local alignment. These methods collectively enhance both the accuracy and efficiency of sequence analysis.
Sequence-to-sequence alignment, naturally extending from classical sequence alignment, has been extensively studied since the 1970s. Dynamic programming methods [29] were among the first to address this problem. Myers and Miller [21] proposed a multiple-sequence alignment algorithm based on the line-sweep paradigm, leveraging range trees for orthogonal range searching to identify the highest-scoring alignment, though with a time complexity increased by a logarithmic factor. In 2005, Abouelhoda and Ohlebusch [1] introduced improved line-sweep algorithms that efficiently solve both global and local fragment-chaining problems in multiple genomes, reducing both time and space complexity compared to earlier approaches. Recent years have seen enhanced models incorporating gaps and overlaps in sequence-to-sequence alignment [4, 18, 12].
To effectively handle the immense volume of genetic sequences and their variations, researchers have increasingly adopted graph-based models, pioneering innovative methods for aligning sequences to a pan-genome graph, which has proven particularly practical [7]. A pan-genome encompasses all genes within a species, composed of a core genome—shared across all individuals—and variable ”dispensable” genomes. Pan-genome graphs, built using graph models, capture the full spectrum of genetic diversity among input genomes. In these graphs, each vertex represents a DNA segment of arbitrary length, while paths correspond to genomes, haplotypes, alleles, or variants. With the aid of a species’ pan-genome graph, sequence alignment aims to find the optimal path that best matches the query sequence, evaluated by alignment score or genomic coverage.
In recent years, both exact and approximation algorithms for sequence-to-graph alignment have received significant attention. When the alignment path is constrained to visit distinct vertices, the problem becomes NP-hard [17] on general pangenome graphs, including De Bruijn graphs [11], as it encompasses the classical Hamiltonian Path problem. However, most existing approaches relax this constraint. Sequence-to-graph alignment on general graphs was initially explored two decades ago in the context of hypertext search [3], where Amir and Lewenstein proposed an $O(|V|+m|E|)$-time algorithm for exact alignment, with $m$ denoting the pattern length. For restricted graph classes, efficient solutions exist: polynomial-time algorithms apply to DAGs [24] and trees [2], with time complexities of $O(|V|+m|E|)$ and $O(|V|)$, respectively. Recently, Equi et al. [9] established a conditional lower bound, demonstrating that, for any constant $\gamma>0$, unless the Strong Exponential Time Hypothesis (SETH) fails, no algorithm can solve the exact alignment problem in $O(|E|^{1-\gamma}m)$ or $O(|E|m^{1-\gamma})$ time on general graphs—even over a binary alphabet.
The approximation scenario is somewhat more intricate than the exact one. When errors are permitted on the query sequence rather than on the graph structure, the problem becomes polynomial-time solvable—not only on DAGs [20] but also on general graphs [3]. In contrast, it remains NP-hard even over a binary alphabet [3, 14]. For De Bruijn graphs, Gibney et al. demonstrated that no sub-quadratic time algorithm exists when substitutions are restricted to the pattern alone [10].
The sub-quadratic term in time complexity poses a significant obstacle, making it practically challenging to process high-throughput sequencing data within a graph-based framework. Recently, the seed-chain-extension approach has become widely adopted in modern alignment tools [13, 15, 26], where co-linear chaining emerges as a key computational challenge.
Co-linear chaining is a well-established concept, extensively studied in sequence-to-sequence alignment [21, 23, 12]. Beyond the query string and reference sequence, a co-linear chaining problem includes a set of anchors—each representing a pair of intervals in the two sequences that correspond to exact seed matches. A valid anchor-chain is an ordered subset of anchors, requiring their intervals to appear in strictly increasing order in both sequences. The goal is to find the highest-scoring chain, where the score combines anchor weights and penalizes gaps between consecutive anchors. Scoring schemes typically incorporate gap costs [23] and may account for overlaps [12].
Nevertheless, co-linear chaining in sequence-to-graph alignment remains more complex than in sequence-to-sequence alignment, posing greater computational challenges. Mäkinen et al. [19] investigated co-linear chaining on directed acyclic graphs (DAGs), developing a sparse dynamic programming approach tailored to DAGs. Subsequently, gap costs were incorporated into the scoring function [6]. More recently, Rajput et al. [25] proposed a practical formulation and an efficient algorithm for co-linear chaining on general graphs, extending gap cost considerations to both query sequences and pan-genome graphs.
In this paper, we investigate the computational complexity of the Gap-sensitive Co-Linear Chaining (Gap-CLC) and Co-Linear Chaining with Errors based on Edit distance (Edit-CLC) in sequence-to-graph alignment. Anchors and anchor-chains in this context are more intricate than their sequence-to-sequence counterparts. We formalize an anchor using the Cartesian Product, as defined in Section
  Notations  , where an anchor is represented not as an explicit tuple, but as an element of a Cartesian Product derived from the occurrences of a sub-sequence on both the query sequence and the pan-genome graph. Consequently, an instance of either the Gap-CLC or Edit-CLC problem comprises a set of sub-sequences along with two associated sets capturing their occurrences on the query and graph. These are formally stated as Problems
  Gap-CLC-form  and
  Edit-CLC-form  , respectively. Section
  complexity  analyzes the computational complexity of both problems. Intuitively, models ignoring ”co-linear” structures can be reduced to the Gap-CLC and Edit-CLC frameworks, implying that incorporating co-linearity does not reduce complexity and may in fact preserve or increase it. We conclude with a summary and discussion of future directions in Section
  conclude  .
2. Basic terminologies and problem definitions
This section begins by formally defining anchor and anchor-chain through the Cartesian Product. It then proceeds to characterize the sequence encoded by an anchor-chain within the context of sequence-to-graph alignment. Finally, it presents the formal formulations of the Gap-CLC and Edit-CLC problems.
A pan-genome graph $G=(V,E,\delta)$ is a directed graph with vertex labels, where $V$ denotes the set of vertices, $E$ denotes the set of directed edges, and $\delta:V\to\Sigma^{*}\setminus\{\epsilon\}$ assigns to each vertex a non-empty string from the finite alphabet $\Sigma$, with $\epsilon$ representing the empty string. We denote the vertex and edge sets of $G$ as $V(G)$ and $E(G)$, respectively. For any vertex $v\in V(G)$, the associated string is given by $\delta(v)$. As reads may span multiple consecutive positions in the graph, we permit self-loops, i.e., an edge $(v,v)\in E(G)$ may exist for some vertex $v$ in $G$.
The number of elements in a sequence $S$ is referred to as its length, denoted by $|S|$. Indices, represented as subscripts, indicate the positions of elements within the sequence, and in this paper, all indexing starts at zero. For instance, in the sequence S=ababcS="ababc", the leftmost ’$a$’ is indexed as $S_{0}$, and the rightmost ’$a$’ is indexed as $S_{2}$. Furthermore, if there exists a walk from vertex $u$ to vertex $v$ in a graph $G$, we say that ”$u$ reaches vv” or equivalently that ”$v$ is reachable from uu”.
Given two non-empty sequences $q\in\Sigma^{*}\setminus\{\epsilon\}$ and $s\in\Sigma^{*}\setminus\{\epsilon\}$, an integer $x$ with $0\leq x<|s|$ is called a location of $q$ in $s$ if $q_{i}=s_{x+i}$ for all $0\leq i<|q|$, provided $x+i<|s|$. In this case, we say $q$ is located at position $x$ in $s$. Clearly, $q$ may occur at multiple locations within $s$. For instance, when $q=ab$ and $s=ababc$, $q$ appears at positions 0 and $2$. Let $\mathrm{loc}(q,s)$ denote the collection of all positions where $q$ appears in $s$.
Given a non-empty sequence $q\in\Sigma^{*}\setminus\{\epsilon\}$ and a pan-genome graph $G$, if there is a vertex $v\in V(G)$ and an integer $x:\ 0\leq x<|\delta(v)|$ such that $q$ is located at $x$ in $\delta(v)$, then the ordered pair $(v,x)$ is said to be a location of $q$ in $G$. And we say $q$ is located at $(v,x)$ in $G$. Let $loc(q,G)$ be the set of all locations of $q$ in the pan-genome graph $G$. For every location $p\in loc(q,G)$, the vertex component is denoted $p\cdot v$ and the offset component $p\cdot x$.
Given two non-empty sequences $q,Q\in\Sigma^{*}\setminus\{\epsilon\}$ and a pan-genome graph $G$, define $A_{q}$ as the cartesian product of $loc(q,Q)$ and $loc(q,G)$, that is, $A_{q}=loc(q,Q)\times loc(q,G)=\{(x,y)\mid x\in loc(q,Q),\,y\in loc(q,G)\}$. It follows that $A_{q}=\emptyset$ if either $loc(q,Q)=\emptyset$ or $loc(q,G)=\emptyset$. Each element in $\bigcup_{q\in\Sigma^{*}\setminus\{\epsilon\}}A_{q}$ is called an anchor, where every anchor is uniquely associated with a sequence $q\in\Sigma^{*}\setminus\{\epsilon\}$. For any anchor $a=(x,y)$, denote its first component as $F(a)=x$ and its second as $H(a)=y$. A subset of $\bigcup_{q\in\Sigma^{*}\setminus\{\epsilon\}}A_{q}$ is referred to as an anchor-set on $(Q,G)$.
Let $Q\in\Sigma^{*}\setminus\{\epsilon\}$ be a non-empty sequence to be aligned with a pan-genome graph $G$, and let $A$ be an anchor-set defined on $(Q,G)$. A pre-order relation $\precsim_{q}$ is established among anchors in $A$ with respect to $Q$: an anchor $a_{i}\in A$ precedes $a_{j}\in A$, denoted $a_{i}\precsim_{q}a_{j}$, if and only if $F(a_{i})<F(a_{j})$. Similarly, with respect to $G$, a pre-order relation $\precsim_{g}$ exists on $A$. We say $a_{i}\precsim_{g}a_{j}$ if either: (1) $H(a_{i})\cdot v\neq H(a_{j})\cdot v$ and there exists a path from $H(a_{i})\cdot v$ to $H(a_{j})\cdot v$ in $G$, or (2) $H(a_{i})\cdot v=H(a_{j})\cdot v$ and $H(a_{i})\cdot x<H(a_{j})\cdot x$. These relations induce a hierarchical ordering of anchors reflecting their sequential and graphical consistency.
For two anchors $a_{i}$ and $a_{j}$, we write $a_{i}\precsim a_{j}$ if both $a_{i}\precsim_{q}a_{j}$ and $a_{i}\precsim_{g}a_{j}$ hold. A sequence of anchors $S$ is called an anchor-chain if and only if $S_{i}\precsim S_{i+1}$ for all $0\leq i<|S|$.
Let $Q\in\Sigma^{*}\setminus\{\epsilon\}$ and $G$ denote a sequence and a pan-genome graph, respectively, and let $W$ be a walk in $G$. The sequence defined by $\sum_{i=0}^{|W|-1}\delta(W_{i})=\delta(W_{0})+\delta(W_{1})+\cdots+\delta(W_{|W|-1})$ is said to be spelled by $W$ in the pan-genome graph, where the operation ”++” represents string concatenation. Notably, every walk in a pan-genome graph inherently corresponds to a spelled sequence.
Recall the Sequence-to-Graph Exact Matching (Exa-SGM) Problem: it seeks a walk in a graph such that the sequence spelled by the vertices along this walk exactly matches a given query sequence. In contrast, the Sequence-to-Graph Matching with Errors (Err-SGM) Problem extends Exa-SGM by aiming to find a walk whose spelled sequence minimizes the edit distance to the query sequence. When each vertex is labeled with a single character, these variants are termed Single-Exa-SGM and Single-Err-SGM, respectively.
2.1 Problem formulations
Besides a query sequence $Q\in\Sigma^{*}\setminus\{\epsilon\}$, an instance of the sequence-to-graph alignment problem—specifically, the Co-Linear Chaining (CLC) problem in sequence-to-graph alignment requires a pan-genome graph $G=(V,E,\delta)$ and an additional anchor-set. Prior studies, such as [21, 23, 12], explicitly define the anchor-set as a collection of tuples. In contrast, this work introduces an implicit definition of the anchor-set through the Cartesian Product of occurrences of certain sequences in $Q$ and $G$, respectively.
An anchor-set $A$ can be decomposed into $k$ disjoint subsets, where each subset contains anchors associated with an identical sequence. Specifically, we write $A=\bigcup_{i=1}^{k}A_{i}$, such that all anchors in $A_{i}$ correspond to the same sequence $q_{i}$, and $q_{i}\neq q_{j}$ whenever $i\neq j$. Thus, the set $R=\bigcup_{i=1}^{k}\{q_{i}\}$ collects all distinct sequences induced by the partition of $A$.
Lemma 1.
Given an alphabet $\Sigma$, a query sequence $Q\in\Sigma^{*}\setminus\{\epsilon\}$, a pan-genome graph $G=(V,E,\delta)$ and a set of anchors $A$ on $(Q,G)$, there are two additional sets $X_{q}\subseteq loc(q,Q)$ and $Y_{q}\subseteq loc(q,G)$ for each sequence $q\in R$, such that $A\subseteq\bigcup_{q\in R}X_{q}\times Y_{q}$.
Proof  Suppose that $A$ can be partitioned into $k$ subsets, i.e., $A=\bigcup_{i=1}^{k}A_{i}$, such that the sequence corresponding to every anchor in $A_{i}$ is $q_{i}$ for $1\leq i\leq k$. According to the definitions of anchors, we have $A_{i}\subseteq A_{q_{i}}=loc(q_{i},Q)\times loc(q_{i},G)$ for $1\leq i\leq k$. Then there are two sets $X_{q_{i}}\subseteq loc(q_{i},Q)$ and $Y_{q_{i}}\subseteq loc(q_{i},G)$ such that $A_{i}\subseteq X_{q_{i}}\times Y_{q_{i}}$ for $1\leq i\leq k$. Therefore, we have $A\subseteq\bigcup_{i=1}^{k}X_{q_{i}}\times Y_{q_{i}}=\bigcup_{q\in R}X_{q}\times Y_{q}$.
By Lemma
  implicit-is-possible  , each explicit anchor-set is a subset of $\bigcup_{q\in R}X_{q}\times Y_{q}$, where $R$ is the set of sequences in the given anchor-set, $X_{q}\subseteq loc(q,Q)$ and $Y_{q}\subseteq loc(q,G)$.
The Gap-CLC problem addressed in this work involves a query sequence $Q\in\Sigma^{*}\setminus\{\epsilon\}$, a pan-genome graph $G=(V,E,\delta)$, and a set of sequences $R\subseteq\Sigma^{*}$. For each sequence $q\in R$, we are given two additional sets: $X_{q}\subseteq loc(q,Q)$ and $Y_{q}\subseteq loc(q,G)$, along with a binary gap-function $f(Gap_{Q}(S),Gap_{G}(S))$. The objective is to find an anchor-chain $S$ from $\bigcup_{q\in R}X_{q}\times Y_{q}$ that maximizes $|S|-f(Gap_{Q}(S),Gap_{G}(S))$, where $Gap_{Q}(S)$ and $Gap_{G}(S)$ denote unary gap-cost functions with respect to $Q$ and $G$, respectively. See Problem
  Gap-CLC-form  for the formal formulation of the Gap-CLC problem. When each vertex in the graph is labeled with a single character, the problem is referred to as the Single-Gap-CLC problem.
Instance: An alphabet $\Sigma$; a query sequence $Q$ over $\Sigma$; a pan-genome graph
$G=(V,E,\delta)$ where $\delta:V\rightarrow\Sigma$; a set of sequences $R\subseteq\Sigma^{*}$; for
each sequence $q\in R$ there are two additionally sets: $X_{q}\subseteq loc(q,Q)$
and $Y_{q}\subseteq loc(q,G)$; a binary gap-function $f(Gap_{Q}(S),Gap_{G}(S))$.
Query: Find an anchor-chain $S$ in $\bigcup_{q\in R}X_{q}\times Y_{q}$ such that
$|S|-f(Gap_{Q}(S),Gap_{G}(S))$ is the maximum.
Problem 1: Gap-sensitive Co-linear Chaining problem (Gap-CLC problem)
The Edit-CLC problem is defined as follows: given a query sequence $Q\in\Sigma^{*}\setminus\{\epsilon\}$, a pan-genome graph $G=(V,E,\delta)$, and a set of sequences $R\subseteq\Sigma^{*}$, along with two associated sets $X_{q}\subseteq loc(q,Q)$ and $Y_{q}\subseteq loc(q,G)$ for each $q\in R$, the goal is to find an anchor-chain $S$ drawn from $\bigcup_{q\in R}X_{q}\times Y_{q}$ such that the edit distance between $Q$ and the sequence induced by $S$ is minimized. Problem
  Edit-CLC-form  proposes a precise formulation of the Edit-CLC problem. When every vertex in the graph is labeled with a single character, this special instance of the Edit-CLC problem is referred to as the Single-Edit-CLC problem.
Instance: An alphabet $\Sigma$; a query sequence $Q$ over $\Sigma$; a pan-genome graph
$G=(V,E,\delta)$ where $\delta:V\rightarrow\Sigma$; a set of sequences $R\subseteq\Sigma^{*}$; for
each sequence $q\in R$ there are two additionally sets: $X_{q}\subseteq loc(q,Q)$
and $Y_{q}\subseteq loc(q,G)$.
Query: Find an anchor-chain $S$ in $\bigcup_{q\in R}X_{q}\times Y_{q}$ such that the
distance between $Q$ and the sequence spelled by $S$ is minimum.
Problem 2: Co-linear Chaining with errors based on Edit distance (Edit-CLC problem)
3. The complexity
In this section, each vertex in the graph is assigned a single-character label, which is sufficient to demonstrate the inherent difficulty of the Gap-CLC and Edit-CLC problems. We present a linear-time reduction from the Single-Exa-SGM problem to Gap-CLC, establishing that no sub-quadratic time algorithm exists for Gap-CLC, unless the Single-Exa-SGM problem admits such an algorithm, which is ruled out even over a binary alphabet [9]. Specifically, for any constant $\gamma>0$, the Single-Exa-SGM problem cannot be solved in either $O(|E|^{1-\gamma}m)$ or $O(|E|m^{1-\gamma})$ time unless the Strong Exponential Time Hypothesis (SETH) fails, where $E$ denotes the edge set of the graph and $m$ is the length of the query sequence [9].
The Single-Exa-SGM problem takes as input a non-empty query string $Q\in\Sigma^{*}\setminus\{\epsilon\}$ and a vertex-labeled graph $G$, with each vertex assigned a label from the binary alphabet $\Sigma=\{0,1\}$. Let $|Q|$ denote the length of $Q$ and $|V(G)|$ represent the number of vertices in $G$. The gap-function $f(Gap_{Q},Gap_{G})$ is defined as follows.
  1. 1.
    The gap function $f(Gap_{Q}(S),Gap_{G}(S))$ is defined as the sum $Gap_{Q}(S)+Gap_{G}(S)$, capturing the combined discrepancy between the two gap measures;
  2. 2.
    The two gap measures are respectively defined as: $Gap_{Q}(S)=\Sigma_{0\leq i<|S|-1}gap_{Q}(S_{i},S_{i+1})$ and $Gap_{G}(S)=\Sigma_{0\leq i<|S|-1}gap_{G}(S_{i},S_{i+1})$;
  3. 3.
    The gap cost associated with $Q$, defined as $\text{gap}_{Q}(S_{i},S_{i+1})=F(S_{i+1})-F(S_{i})-1$, is computed for each consecutive pair $(S_{i},S_{i+1})$ with $0\leq i<|S|-1$, capturing the incremental difference in the function $F$ minus a unit offset;
  4. 4.
    For each $0\leq i<|S|-1$, the gap cost associated with $G$ is defined as $gap_{G}(S_{i},S_{i+1})=0$ if $H(S_{i+1})\cdot v$ is adjacent to $H(S_{i})\cdot v$ in the graph; otherwise, $gap_{G}(S_{i},S_{i+1})=|Q|+|V(G)|$,
where $S$ is an anchor-chain.
The reduction from Single-Exa-SGM to Gap-CLC proceeds as follows.
  1. 1.
    The alphabet continues to be $\{0,1\}$, and likewise, the query sequence and the pan-genome graph remain denoted as $Q$ and $G$, respectively;
  2. 2.
    The set of sequences is $R=\{0,1\}$;
  3. 3.
    Let $X_{q}=loc(q,Q)$ and $Y_{q}=loc(q,G)$ for every $q\in R$;
It is easy to find that the reduction takes linear-time. Recall that the sequence adhering to every anchor is a single character. So both $loc(q,Q)$ and $loc(q,G)$ for $q\in R$ can be obtained by traversing $Q$ and $G$ in linear time.
Lemma 2.
The reduction takes $O(|Q|+|V(G)|)$-time.
Proof  The alphabet, query sequence, pan-genome graph, and the set of sequences $R$ are identical to those defined in the Single-Exa-SGM problem, with $R$ sharing the same composition as the alphabet. Consequently, the construction of these components—alphabet, query sequence, pan-genome graph, and RR—can be accomplished in constant time.
Moreover, obtaining the location functions, $loc(q,Q)$ and $loc(q,G)$, requires $O(|Q|+|V(G)|)$ time, and defining the binary gap-function also takes $O(|Q|+|V(G)|)$ time. Thus, the entire reduction runs in $O(|Q|+|V(G)|)$ time.
Let $\Pi$ be the set of instances obtained by the reduction. Let $\omega$ == $|S|$ - $($$Gap_{Q}(S)$ ++ $Gap_{G}(S)$)) be the objective function of the Single-Gap-CLC problem, where $S$ is an anchor-chain.
Lemma 3.
For each instance in $\Pi$, if there is an anchor-chain $S$ in $\bigcup_{q\in R}loc(q,Q)\times loc(q,G)$ such that $\omega\geq|Q|$, then the number of anchors in $S$ is exact $|Q|$.
Proof  Since $\omega\geq|Q|$, the set $S$ contains at least $|Q|$ anchors. We now demonstrate that the number of anchors in $S$ cannot exceed $|Q|$, establishing the exact count.
Suppose, to the contrary, that $|S|>|Q|$. By definition, for any two anchors $S_{i}$ and $S_{j}$ in $S$ with $i<j$, we have $F(S_{i})<F(S_{j})$. Since $Q$ contains exactly $|Q|$ elements, there are only $|Q|$ distinct positions available in $Q$. If $S$ had more than $|Q|$ anchors, then by the pigeonhole principle, at least two distinct anchors $S_{i}$ and $S_{j}$ in $S$ would occupy the same position in $Q$, implying $F(S_{i})=F(S_{j})$. This contradicts the strict monotonicity required by the anchor-chain definition. Hence, $|S|$ cannot exceed $|Q|$, and the number of anchors in $S$ must be exactly $|Q|$.
Lemma 4.
For each instance in $\Pi$, there is an anchor-chain $S$ in $\bigcup_{q\in R}loc(q,Q)\times loc(q,G)$ such that $\omega\geq|Q|$, if and only if there is a walk $P$ in $G$ such that the sequence spelled by $P$ is the same as $Q$.
Proof  Suppose that there is an anchor-chain $S$ such that $\omega\geq|Q|$. According to Lemma
  anchor-nu  , $S$ contains precisely $|Q|$ anchors in $S$. This implies that every character of $Q$ is included in the resulting solution. Furthermore, both $Gap_{Q}(S)=0$ and $Gap_{G}(S)=0$ must hold simultaneously, indicating that $F(S_{i+1})=F(S_{i})+1$ for any consecutive anchors $S_{i}$ and $S_{i+1}$ in $S$. Additionally, $H(S_{i})\cdot v$ and $H(S_{i+1})\cdot v$ are adjacent in $G$ for all $0\leq i<|S|-1$. Hence, $S$ forms a walk in $G$ where the sequence induced by $S$ exactly spells $Q$.
Consider a walk $P$ on $G$ whose vertex sequence spells out the same string as $Q$. Since each vertex in $G$ is labeled with a single character, the length of $P$ (in terms of vertices) exactly matches the number of characters in $Q$. Let $S$ be an anchor-chain with $|S|=|P|$, where $F(S_{i})=i$ and $H(S_{i})=(P_{i},0)$ for $0\leq i\leq|P|-1$. Then, $Gap_{Q}(S)=0$, $Gap_{G}(S)=0$, and $|S|=|Q|$. Hence, we obtain an anchor-chain $S$ in $\bigcup_{q\in R}loc(q,Q)\times loc(q,G)$ such that $\omega\geq|Q|$.
Theorem 1.
There is no sub-quadratic time algorithm for the Gap-CLC problem, unless the Strong Exponential Time Hypothesis (SETH) is false, even the alphabet is binary.
Proof  Suppose, to the contrary, that there exists a sub-quadratic time algorithm $\mathcal{F}$ for the Gap-CLC problem. Then, $\mathcal{F}$ can solve the Single-Gap-CLC problem on the instance set $\Pi$ in sub-quadratic time.
By Lemma
  one-to-one-corr-Gap  , for each instance in $\Pi$, whenever the objective value $\omega$ satisfies $\omega\geq|Q|$, there exists a walk in the pan-genome graph whose induced sequence exactly matches the given query sequence $Q$. Conversely, if $\omega<|Q|$, no such walk can spell the entire sequence $Q$. Furthermore, according to Lemma
  reduction-complexity  , constructing the instances in $\Pi$ requires sub-quadratic time.
Consequently, we have developed a sub-quadratic time algorithm for the Single-Exa-SGM problem, challenging the recent assertion that no such algorithm exists unless the Strong Exponential Time Hypothesis (SETH) fails—even over a binary alphabet [9]. This implies that the Gap-CLC problem also admits no sub-quadratic time algorithm.
Moreover, the Single-Err-SGM problem has been shown to be NP-hard when errors are permitted on a graph [3], even in the case of a binary alphabet [14]. We present a linear-time reduction from the Single-Err-SGM problem to the Edit-CLC problem and establish its NP-hardness in Theorem
  Err-CLC-NP-hard  .
Theorem 2.
The Edit-CLC problem is NP-hard if errors are allowed on pan-genome graphs, even on a binary alphabet.
Proof  We further augment the Single-Err-SGM problem instance by appending the sequence set $R=\{0,1\}$, and define $X_{q}=\text{loc}(q,Q)$ and $Y_{q}=\text{loc}(q,G)$ for each $q\in R$. This transformation effectively reduces the Single-Err-SGM problem to the Single-Edit-CLC problem.
In addition, this reduction can be computed in $O(|Q|+|V(G)|)$-time. Recall that Amir et al. [3] demonstrated the NP-hardness of the Single-Err-SGM problem when errors are permitted in the graph, even over a binary alphabet [14]. Consequently, the Edit-CLC problem is also NP-hard when errors are allowed in pan-genome graphs.
4. Conclusions and discussions
Co-linear chaining problems have found extensive applications in sequence-to-sequence alignment, yet their extension to sequence-to-graph alignment presents greater challenges. In this work, we investigate the computational complexity of co-linear chaining in the context of sequence-to-graph alignment, focusing specifically on the Gap-CLC and Edit-CLC problems. We demonstrate that the Gap-CLC problem is at least as hard as the Exa-SGM problem, while the Edit-CLC problem is proven to be NP-hard when errors are permitted on pan-genome graphs—even over a binary alphabet. These complexity results reveal that incorporating the ”co-linear” constraint in sequence-to-graph alignment does not reduce the theoretical complexity compared to non-co-linear models. Despite these theoretical challenges, ”co-linear” chaining remains widely adopted in practical tools, highlighting a significant gap between theoretical understanding and real-world practice. Bridging this gap presents an important and intriguing direction for future research. Furthermore, algorithmic solutions for the Edit-CLC problem remain scarce, with limited progress on both exact and approximation algorithms, indicating a clear need for further investigation and development in this area.
Moreover, in this study, we do not define an anchor as an explicit tuple in the formulations of the Gap-CLC and Edit-CLC problems. While Lemma
  implicit-is-possible  indicates that every explicit anchor-set is contained within $\bigcup_{q\in R}X_{q}\times Y_{q}$, where $R$ denotes the set of sequences in the anchor-set and $X_{q}\subseteq loc(q,Q)$, $Y_{q}\subseteq loc(q,G)$, adopting implicit anchor-sets alters the input structure of co-linear chaining problems. It remains an open question whether a linear-time or even a sub-quadratic time reduction exists when anchors are represented as explicit tuples.